OpenStudy (anonymous):

Identify the center and radius for the following equation: x²+y²-6x+10y-15=0

OpenStudy (raden):

the general form an equation of circle is : x^2 + y^2 + Ax + By + C = 0, with A, B, and C are the constants so to get the center of circle and its radius, u can use the formula : the center is (-A/2 , -B/2) r = sqrt(A^2/4 + B^2/4 - C)

OpenStudy (anonymous):

thank you!! is that the final answer?

OpenStudy (raden):

that is the formula only... so, if given the equation : x²+y²-6x+10y-15=0, it means A=-6, B=10, C=-15 now, apply that numbers to formula above ^, what u get

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

thanks!! @RadEn so what would the center and radius be from that?

OpenStudy (raden):

let the coordinat of center is (h, k) with, h=-A/2 = -(-6)/2 = 6/2 = .... k = -B/2 = -10/2 = ...... so, the coordinat center is (... , ...) r = sqrt (A^2/4 + B^2/4 - C) r = sqrt[ (-6)^2/4 + (10)^2/4 - (-15)] r = sqrt(36/4 + 100/4 + 15) r = sqrt(9 + 25 + 15) r = sqrt(49) = ....

OpenStudy (anonymous):

so r=7 and center= (3,-5)?? thank u so much :) @raden

OpenStudy (raden):

yeah, that's right welcome :*